# Giulia Gaggero and Elisa Gorla

### v2

@@ 4 The degree of regularity of a random system
The next theorem yields an upper bound on the degree of regularity of an algebraically random system of equations of degree larger than the field size to which we add the field equations. For such a system, the bound is tighter than the one from Corollary 28. We start with a preparatory lemma, whose proof follows directly from the definition.
@@
Assume that Conjecture 19 holds. Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, with $q\leq D\leq n(q-1)$. Assume that $m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}$ and that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$. Let $11\leq t\leq q-$ 1 and $11\leq k\leq n-$ 1 be such that $m$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
$\sigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
@@
$B=q-t+(n-k)(q-1).$
If $m=\sigma_{k,t}$ and either $t\neq q-1$ and $D\geq 2q-t-1$, or $t=q-1$ and $D\geq 2q-1$, then
@@
We start by proving that $\eta_{k,t}$ is the number of $(D,k,q-t)$-type monomials in $R$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$, that is
$\eta_{k,t}=\dim\left[\frac{\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}.$ (3)
Since $x_{k+1}^{q},\ldots,x_{n}^{q}$ is a regular sequence in $\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]$, a standard computation involving Hilbert series yields the explicit formula
$\dim\left[\frac{\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}=\sum_{i=0}^{\left\lfloor\frac{D+t}{q}\right\rfloor-1}(-1)^{i}\binom{n-k}{i}\binom{n-k-1+D-(i+1)q+t}{n-k-1},$
@@
$\sigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
is the number of monomials of degree $D$ which do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$ and are greater than or equal to $x_{k}^{q-t}x_{n}^{D-q+t}$, the smallest $(D,k,q-t)$-type monomial. Moreover, $\sigma_{k-1,q-1}=\sigma_{k,0}$ for $12\leq k\leq n-$ 1.
Let $I\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ be the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\}$. Let $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$ be the $(q,\ldots,q;D)$-LPP ideal such that
@@
Hence $\mathcal{L}$ is minimally generated by $x_{1}^{q},\ldots,x_{n}^{q}$ and $m$ monomials of degree $D$ that do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$. Recall that, by assumption, $L$ is the largest lex-segment ideal generated in degree $D$ such that $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$, i.e., $L_{D}\supseteq(x_{1}^{q},\ldots,x_{n}^{q})_{D}$.
Let $u$ be the smallest degree $D$ monomial in $\mathcal{L}/(x_{1}^{q},\ldots,x_{n}^{q})$. Notice that $u$ is a $(D,k,q-t)$-type monomial, since $m$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$. Let $v$ be the smallest degree $D$ monomial in $L$. If $u$ is the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$, then $u=x_{k}^{q-t}x_{k+1}^{q-1}\cdots x_{n}^{q-1}$ and $D=q-t+(n-k)(q-1)$. Moreover, $\mathcal{L}_{D}=R_{D}$ and $v=x_{n}^{D}$ is a $(D,n,D)$-type monomial.
Assume now that $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ and let $w$ be the monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ which follows $u$ in decreasing lexicographic order. Then $u\geq v\geq w$ and $v$ is the degree $D$ monomial next to $w$ in increasing lexicographic order. If $m\neq\sigma_{k,t}$, then $w$ has type $(D,k,q-t)$, hence so does $v$ by Lemma 33. Suppose therefore that $m=\sigma_{k,t}$. Write $D-q+t=(n-\ell)(q-1)+r$, where $10\leq r<q-$ 1. Notice that $D-q+t<(n-k)(q-1)$, since $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$. Then $\ell>k$. In this situation, $u=x_{k}^{q-t}x_{\ell}^{r}x_{\ell+1}^{q-1}\cdots x_{n}^{q-1}$ and $w=x_{k}^{q-t-1}x_{k+1}^{q-1}\cdots x_{k+n-\ell}^{q-1}x_{k+n-\ell+1}^{r+1}$. Notice that $w$ is a $(D,k,q-t-1)$-type monomial if $t\neq q-1$ and a $(D,k+1,q-1)$-type monomial if $t=q-1$. If $w$ is not the smallest degree $D$ monomial of its type in $R_{D}$, then $v$ has the same type as $w$. If $w$ is the smallest degree $D$ monomial of its type in $R_{D}$, then $w=x_{k}^{q-t-1}x_{k+1}^{D-q+t+1}$ in the case $t\neq q-1$ and $w=x_{k+1}^{q-1}x_{k+2}^{D-q+1}$ in the case $t=q-1$. Since $w\not\in(x_{1}^{q},\ldots,x_{n}^{q})$, this is only possible if $D\leq 2q-t-2$ for $t\neq q-1$ and $D\leq 2q-2$ for $t=q-1$. In this situation, $v$ has type $(D,k,q-t)$ in the case $t\neq q-1$ and type $(D,k+1,q)$ in the case $t=q-1$.
@@
$\operatorname{reg}(\mathcal{L})=\left\{\begin{array}[]{ll}q-t-1+(n-k)(q-1)&\mbox{if }m=\sigma_{k,t}\mbox{ and }\\ &\mbox{either }t\neq q-1\mbox{ and }D\geq 2q-t-1,\\ &\mbox{or }t=q-1\mbox{ and }D\geq 2q-1,\\ \\ q-t+(n-k)(q-1)&\mbox{else.}\end{array}\right.$ (5)
@@
5 Applications to the study of GeMSS and Rainbow
GeMSS and Rainbow were the only multivariate schemes in Round 3 of the NIST Post-Quantum Cryptography Standardization process. They are based on modifications of HFE (Hidden Field Equation) and UOV (Unbalanced Oil and Vinegar), respectively. They were subsequently revealed to be insecure and susceptible to a MinRank Attack, see e.g. [3, 8, 9, 1], therefore they were excluded from the NIST competition.
In this section, we demonstrate how to use our results to show that these systems are far from being algebraically random. We do so by computing their degree of regularity or solving degree for small instances and comparing it with the upper bounds that we obtained in Theorem 22 and Theorem 34. The degree of regularity and solving degree of the systems associated to GeMSS and Rainbow are much smaller than the corresponding invariants for an algebraically random system with the same parameters, which reveals the presence of a hidden structure that may be used to mount an ad-hoc attack, as it was done in practice with the MinRank Attacks mentioned above.
We used Magma to compute the solving degree and Singular to compute the degree of regularity. The values that we obtain (and that we indicate in the tables below) are almost always the same for systems with the same parameters. For each choice of the parameters in the table, we produce ten instances of the public key $\mathcal{PK}=\{p_{1},\ldots,p_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of the chosen scheme. For each one of them, we choose a random vector $s=(s_{1},\ldots,s_{m})\in\mathbb{F}_{q}^{m}$ as a signature (in case the chosen vector is not a valid signature, we replace it with another randomly chosen vector). In order to forge the signature, an attacker may want to solve the system $\mathcal{PK}s=\{p_{1}-s_{1},\ldots,p_{m}-s_{m}\}$. We make the system $\mathcal{PK}s$ square by assigning random values to the last $n-m$ variables. This yields a system =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{m}]$.
Since in GeMSS we work over $\mathbb{F}_{2}$, we add to the system $\mathcal{F}$ the field equations =$\mathcal{E}=\{x_{1}^{2}+x_{1},\ldots,x_{m}^{2}+x_{m}\}$. In the next table we compare the experimental results we obtained for GeMSS with the bounds from Theorem 34. The experiments show that both the solving degree and the degree of regularity of $\mathcal{F}$ can be more than twice the solving degree of $\mathcal{F}\cup\mathcal{E}$. This confirms the intuition that adding the field equations is a good strategy in order to solve the system $\mathcal{F}$ over $\mathbb{F}_{2}$.
Unfortunately we were able to compute the degree of regularity of $\mathcal{F}\cup\mathcal{E}$ only for small values of the parameters. In the next table the first three columns contain the parameters of the cryptosystem, and the fourth the number of polynomials and variables that appear in $\mathcal{F}$. The columns labelled '$d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$' and '$\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contain the values computed with Magma and Singular. The columns labelled '$\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' and '$\max d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$' are the bounds given by the Theorem 34 for the chosen parameters and a standard Gröbner basis algorithm.
n, D a v m $d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\max d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$ $\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$
12, 4 1 1 11 5 10 3 18
8, 9 1 1 7 3 6 3 10
8, 9 1 2 7 3 6 3 10
8, 9 2 1 6 3 5 3 8
8, 9 2 2 6 3 5 3 8
24, 4 1 1 23 22 4 42
24, 4 1 2 23 22 4 42
24, 4 1 3 23 22 4 42
24, 4 2 1 22 21 4 40
24, 4 2 2 22 21 4 40
24, 4 3 1 21 20 4 38
While GeMSS is random according to Definition 12, the experimental results make it clear that both the degree or regularity and the solving degree of GeMSS are far from the largest values that one can find for a system of those parameters according to Theorem 34. This indicates that, for the same parameters, one should be able to find systems for which the complexity of computing a Gröbner basis is much larger. More importantly, it reveals the presence of a hidden algebraic structure, which may be exploited in ad-hoc attack (as it was in fact done in the MinRank attacks mentioned in the opening paragraph).
For Rainbow, we choose to work over $\mathbb{F}_{4}$ and $\mathbb{F}_{9}$. Since $\mathcal{F}$ is a square system, $\mathcal{F}$ is algebraically random if and only if $\mathcal{F}^{\operatorname{\mathrm{top}}}$ is a regular sequence. This turns out to be the case in most of the examples that we computed and in that case
$d_{\operatorname{reg}}(\mathcal{F})=m+1$ (6)
by Theorem 22. This is confirmed by our computations.
Since the systems coming from this scheme are quadratic, adding the field equations may increase the solving degree of the system. However, for the small values of $q$ that we tried in our experiments, we find that in all cases but one the solving degree decreases when adding the field equations. This makes sense, as the degree $q$ of the equations that we add is never larger than the solving degree of the system to which we add them. In the next table we summarize the results that we obtained in our computational experiments. Since in our experiments the solving degree of $\mathcal{F}\cup\mathcal{E}$ is almost always smaller than that of $\mathcal{F}$, in our examples $\mathcal{F}\cup\mathcal{E}$ is the relevant system to consider, that is, the system that one wants to try to solve. Therefore, we consider the degree of regularity and solving degree of $\mathcal{F}\cup\mathcal{E}$. In our table, we compare the degree of regularity and the solving degree of $\mathcal{F}\cup\mathcal{E}$ with the upper bounds from Theorem 22 and Corollary 28. We use Corollary 28 as the algorithm implemented in Magma is a standard one. The first three columns contain the chosen values for the parameters and the number of polynomials and variables that appear in $\mathcal{F}$. The columns labelled '$d_{\operatorname{reg}}(\mathcal{F})$' and '$\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contain the values computed with Magma and Singular. The column labelled '$\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$' contains the bounds from Corollary 28. We do not include the values of the degree of regularity and solving degree of $\mathcal{F}$ in the table, as experimentally we find that the system $\mathcal{F}\cup\mathcal{E}$ can always be solved more efficiently than the system $\mathcal{F}$.
$q$ $[v_{1},o_{1},o_{2}]$ $m$ $d_{\operatorname{reg}}(\mathcal{F})$ $d_{\operatorname{reg}}(\mathcal{F}\cup\mathcal{E})$ $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$ $\max\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})$
4 $[3,2,2]$ 4 4/5 4 4 8
4 $[3,3,3]$ 6 6/7 5 5 12
4 $[7,5,5]$ 10 10/11 6 6 20
9 $[3,2,2]$ 4 5 5 9
9 $[7,5,5]$ 10 11 9/10 10 20
Notice that Corollary 28 does not apply to the case $q=9$ and $[v_{1},o_{1},o_{2}]=[3,2,2]$, since $d_{\operatorname{reg}}(\mathcal{F})=5$. For these parameters, the bound from Corollary 28 would yield $\operatorname{\mathrm{solv.deg}}(\mathcal{F}\cup\mathcal{E})\leq 8$. However, the bound does not hold in this case, as our experiments show.
As for GeMSS, we observe that the values that we computed for the solving degree of $\mathcal{F}\cup\mathcal{E}$ are far from the upper bounds predicted by Corollary 28. We conclude that, also in this case, one expects to find systems with the same parameters as these instances of Rainbow and for which the complexity of computing a Gröbner basis is larger. More importantly, once again it reveals the presence of a hidden algebraic structure, which may be exploited in ad-hoc attack (as it was in fact done in the attacks that we mentioned at the start of the section).

### v3

@@ 4 Bounds on the degree of regularity
The next theorem yields an upper bound on the degree of regularity of a system of equations of degree $D$ larger than the field size, to which we add the field equations. We start with a preparatory lemma, whose proof follows directly from the definition.
@@
The next theorem applies, e.g., to a system of equations of the same degree $D\geq 2$ over $\mathbb{F}_{2}$. Notice that, by Lemma 30, the assumption that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$ implies that $m\leq\sum_{i=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{i}\binom{n}{i}\binom{n+D-1-iq}{n-1}$.
@@
Assume that Conjecture 20 holds. Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, with $q\leq D\leq n(q-1)$. Assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$. Let $1\leq t\leq q-1$ and $1\leq k\leq n-1$ be such that $m$ belongs to the interval $(\varsigma_{k,t-1},\varsigma_{k,t}]$, where
$\varsigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
@@
B=qt+(nk)(q1).B=q-t+(n-k)(q-1).
If $m=\varsigma_{k,t}$ and either $t\neq q-1$ and $D\geq 2q-t-1$, or t=q1t=q-1 and $D\geq 2q-1$, then
@@
The idea of the proof is similar to that of Theorem 23. In order to apply Proposition 21, we need to determine the value of $k$ and $t_{k}$ for the monomial in position $m$ in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{q},\ldots,x_{n}^{q})$. As we will show, these values are the unique $k$ and $t_{k}$ such that $m$ belongs to the interval $(\varsigma_{k,q-t_{k}-1},\varsigma_{k,q-t_{k}}]$.
We start by proving that $\eta_{k,t}$ is the number of (D,k,qt)(D,k,q-t)-type monomials in $R$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})_{D}$, that is
$\eta_{k,t}=\dim\left[\frac{\mathbb{F}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}.$ (4)
Since $x_{k+1}^{q},\ldots,x_{n}^{q}$ is a regular sequence in $\mathbb{F}[x_{k+1},\ldots,x_{n}]$, a standard computation involving Hilbert series yields the explicit formula
$\dim\left[\frac{\mathbb{F}[x_{k+1},\ldots,x_{n}]}{(x_{k+1}^{q},\ldots,x_{n}^{q})}\right]_{D-q+t}=\sum_{i=0}^{\left\lfloor\frac{D+t}{q}\right\rfloor-1}(-1)^{i}\binom{n-k}{i}\binom{n-k-1+D-(i+1)q+t}{n-k-1}$,
@@
$\varsigma_{k,t}=\sum_{i=1}^{k-1}\sum_{j=1}^{q-1}\eta_{i,j}+\sum_{j=1}^{t}\eta_{k,j}$
is the number of monomials of degree $D$ which do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$ and are greater than or equal to $x_{k}^{q-t}x_{n}^{D-q+t}$, the smallest (D,k,qt)(D,k,q-t)-type monomial. Moreover, $\varsigma_{k-1,q-1}=\varsigma_{k,0}$ for $2\leq k\leq n-1$.
Let $I\subseteq\mathbb{F}[x_{1},\ldots,x_{n}]$ be the ideal generated by $\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\}$. Let $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$ be the $(q,\ldots,q;D)$-LPP ideal such that
@@
Hence $\mathcal{L}$ is minimally generated by $x_{1}^{q},\ldots,x_{n}^{q}$ and $m$ monomials of degree $D$ that do not belong to $(x_{1}^{q},\ldots,x_{n}^{q})$. Recall that, by assumption, $L$ is the largest lex-segment ideal generated in degree $D$ such that $\mathcal{L}=(x_{1}^{q},\ldots,x_{n}^{q})+L$.
Let $u$ be the smallest degree $D$ monomial in $\mathcal{L}/(x_{1}^{q},\ldots,x_{n}^{q})$. Notice that $u$ is a (D,k,qt)(D,k,q-t)-type monomial, since $m$ belongs to the interval $(\varsigma_{k,t-1},\varsigma_{k,t}]$. Let $v$ be the smallest degree $D$ monomial in $L$. If $u$ is the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$, then $u=x_{k}^{q-t}x_{k+1}^{q-1}\cdots x_{n}^{q-1}$ and D=qt+(nk)(q1)D=q-t+(n-k)(q-1). Moreover, $\mathcal{L}_{D}=R_{D}$ and $v=x_{n}^{D}$ is a $(D,n,D)$-type monomial.
Assume now that $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ and let $w$ be the monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ which follows $u$ in decreasing lexicographic order. Then $u\geq v\geq w$ and $v$ is the degree $D$ monomial next to $w$ in increasing lexicographic order. If $m\neq\varsigma_{k,t}$, then $w$ has type (D,k,qt)(D,k,q-t), hence so does $v$ by Lemma 32. Suppose therefore that $m=\varsigma_{k,t}$. Write $D-q+t=(n-\ell)(q-1)+r$, where $0\leq r<q-1$. Notice that Dq+t<(nk)(q1)D-q+t<(n-k)(q-1), since $u$ is not the smallest monomial in $(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$. Then $\ell>k$. In this situation, $u=x_{k}^{q-t}x_{\ell}^{r}x_{\ell+1}^{q-1}\cdots x_{n}^{q-1}$ and $w=x_{k}^{q-t-1}x_{k+1}^{q-1}\cdots x_{k+n-\ell}^{q-1}x_{k+n-\ell+1}^{r+1}$. Notice that $w$ is a (D,k,qt1)(D,k,q-t-1)-type monomial if $t\neq q-1$ and a (D,k+1,q1)(D,k+1,q-1)-type monomial if t=q1t=q-1. If $w$ is not the smallest degree $D$ monomial of its type in $R_{D}$, then $v$ has the same type as $w$. If $w$ is the smallest degree $D$ monomial of its type in $R_{D}$, then $w=x_{k}^{q-t-1}x_{k+1}^{D-q+t+1}$ in the case $t\neq q-1$ and $w=x_{k+1}^{q-1}x_{k+2}^{D-q+1}$ in the case t=q1t=q-1. Since $w\not\in(x_{1}^{q},\ldots,x_{n}^{q})$, this is only possible if $D\leq 2q-t-2$ for $t\neq q-1$ and $D\leq 2q-2$ for t=q1t=q-1. In this situation, $v$ has type (D,k,qt)(D,k,q-t) in the case $t\neq q-1$ and type $(D,k+1,q)$ in the case t=q1t=q-1.
@@
[…]
@@
We state as a corollary the case of a system of binary equations. This is a case of special interest and has a simpler formulation than the general case.
Corollary 34.
Assume that Conjecture 20 holds. Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a system of degree $D$ polynomials in $n$ variables over $\mathbb{F}_{2}$, with $2\leq D\leq n$. Assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $x_{1}^{2},\ldots,x_{n}^{2}$. Let $1\leq k\leq n-1$ be such that $m$ belongs to the interval $(\varsigma_{k-1},\varsigma_{k}]$, where
$\varsigma_{k}=\sum_{i=1}^{k}{n-i\choose D-1}$.
If $m=\varsigma_{k}$ and and $D\geq 3$, then
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-k,$
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq 2(n-k-1),$
and
$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-k+1.$
Else
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-k+1,$
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq 2(n-k),$
and
$\operatorname{\mathrm{solv.deg}}^{m}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq n-k+2.$
Corollary 33 applies to, e.g., the Unbalanced Oil and Vinegar scheme (UOV) from [32, Section 2], which over $\mathbb{F}_{2}$ is a system of $m$ quadratic equations in $n>m$ variables, to which one adds the $n$ quadratic field equations. It also applies to the Oil and Vinegar scheme of degree three from [32, Section 9] and to the Unbalanced Oil, Vinegar and Salt scheme from [32, Section 11]. Over $\mathbb{F}_{2}$, the first is a system of $m$ cubics in $n>m$ variables, to which one adds the $n$ quadratic field equations. The second is a system of $m$ quadratic equations with the Oil and Vinegar structure in $n>m$ variables, to which one adds $r$ random quadratic equations and the $n$ quadratic field equations.
Example 35.
Let $\mathcal{F}$ be an UOV system in degree $2\leq D\leq n$ over $\mathbb{F}_{2}$. For a UOV system one always has $m<n$, since nmn-m is the number of vinegar variables. If $D=n$, then $m=\varsigma_{1}=1$. Else, $\varsigma_{1}={n-1\choose D-1}\geq n-1$. In both cases, one has $m\in(0,\varsigma_{1}]$ and $k=1$. Therefore, by Corollary 34
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n$
and if either $D=n$ and $m=1$, or $D\geq 3$ and m=n1m=n-1, then
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{2}-x_{1},\ldots,x_{n}^{2}-x_{n}\})\leq n-1.$
One can also approach a UOV system by assigning random values to nmn-m variables and trying to solve for the rest. This yields a system of $m$ equations in $m$ variables, to which one can still add the field equations. We compute the bound for a quadratic UOV system.
Example 36.
Let $\mathcal{F}$ be an UOV system in degree $D=2$ over $\mathbb{F}_{2}$. Assign random values to nmn-m variables to obtain a system of $m$ quadratic equations in $m$ variables. If $m>2$, then $m\leq\dim_{\mathbb{F}_{2}}\mathbb{F}_{2}[x_{1},\ldots,x_{m}]_{2}-\dim_{\mathbb{F}_{2}}\langle x_{1}^{2},\ldots,x_{m}^{2}\rangle$, so one expects that the specializations of $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ remain linearly independent modulo $\langle x_{1}^{2},\ldots,x_{m}^{2}\rangle$. Since $\varsigma_{1}=m-1$ and $\varsigma_{2}=2m-3$, then $m\in(\varsigma_{1},\varsigma_{2}]$ and $k=2$. Therefore, by Corollary 34 the degree of regularity is bounded from above by m1m-1.
In Theorem 23 and Theorem 33 we consider a system $\mathcal{F}$ of equations of the same degree $D$ and provided upper bounds on the degree of regularity of $\mathcal{F}$ under the assumption that $\mathcal{F}$ is regular in degree $D$, and of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ if $D\geq q$. Since our results also apply to semiregular systems, the bounds that we obtain are necessarily larger than the value of the degree of regularity of a semiregular system of equations of the same degree. A common approach in the cryptoanalysis of a multivariate scheme is to assume that the corresponding system is semiregular. Under such an assumption, one can compute the value of the degree of regularity of the system. We stress that, even though the bounds that we obtain are less tight than those that one obtains by assuming that a system is semiregular, they have the advantage of being provable, since they just require that the degree of regularity is finite and do not rely on the heuristic assumption that the system is semiregular.